OpenStudy (anonymous):

use mathematical induction to prove the formula for every positive integer n. 3+7+11+15+...+(4n -1)=n(2n+1)

OpenStudy (zzr0ck3r):

4n-1 does not equal n(2n+1)

OpenStudy (zzr0ck3r):

for any n in N

OpenStudy (cwrw238):

let n = k then if formula is true: sum of k terms = k(2k + 1) then sum of k+1 terms = k(2k + 1) + (4(k+1) - 1 = 2k^2 + k + 4k + 4 - 1 = 2k^2 + 5k + 3 = (2k + 3 )(k + 1 ) or (k + 1)(2(k + 1) + 1) which has the same form as the formula for sum of k terms so if its true for k terms then it is also true for k + 1 terms it is true for first term so it must be true for 2nd term , then true for 3rd etc so its true for all values of n

OpenStudy (zzr0ck3r):

you sure you wrote ht eproblem right?

OpenStudy (cwrw238):

* (k + 1)(2(k + 1) + 1) has same form as k(2k + 1) with k being replaced by (k+1)

OpenStudy (cwrw238):

do you follow that garlucia?

OpenStudy (zzr0ck3r):

so im confused but this does not seem to be true for any n, so what is the point of induction?

OpenStudy (cwrw238):

well its true for 1 term: = 1(2*1 + 1) = 3 and if its true for kWh term it is also true for (k+1)th term because (k+1) replaces k in the formula so true for term then it must be true for term , term 3, 5 and so on thats method of induction

OpenStudy (cwrw238):

* true for term 1 , must be true for 2 , then true for term 3 and so on

OpenStudy (zzr0ck3r):

but its not true for n = 3 (4*3-1) = 11 3(2*3+1) = 21

OpenStudy (cwrw238):

sum = n(2n + 1) sum of 3 terms = 3(2(3) + 1) = 3* 7 = 21 - you are using the formula for nth term (4n-1) not formula for the sum of n terms

OpenStudy (zzr0ck3r):

o snap sorry.....its been a long day

OpenStudy (cwrw238):

lol thats ok